# A generalized statistical model for fits to parton distributions

### v1

@@ II.2 Least Squares Fits and Tension
Figure 2: Panel (a): the comparison of pseudo-data and optimal pseudo-PDFs. The uncertainties in replica sets correspond to the 1 standard deviation interval. For the LS fit $C$, the Hessian uncertainty with $\Delta\chi^{2}=1$ is also provided, which deviates from the Monte Carlo uncertainty by a negligible amount. The sizes of uncertainties are small in view of the magnitudes of pseudo-PDFs. Panel (b): similar to panel (a), but for ratios to the central value of the Hessian set of the SG-C fit. Panel (c): clouds of fitted values of parameters $a_{4}$ and $a_{5}$ in replica sets, as compared to true values.
@@ III Gaussian mixture model in PDF fit
As mentioned in the earlier section, the likelihood from each of the data sets (pseudo-data #1 and pseudo-data #2) are significantly different and needs to be combined in a meaningful way. We propose the use of the Gaussian Mixture Model to construct the combined likelihood for data with tension.
The GMM was first used by Karl Pearson in 1894 [30]. It is now used extensively as an unsupervised machine learning technique to classify data [31]. In general, the GMM is a mixture of $K$ base Gaussian distributions. Given the data set $Y=\{(y_{1},\Delta y_{1}),(y_{2},\Delta y_{2}),\cdots,(y_{N_{\text{pt}}},\Delta y_{N_{\text{pt}}})\}$, where $\Delta y_{j}$ is the uncertainty associated with the measured value $y_{j}$222 For clarity of presentation, we use an uncorrelated data set and present formulas for this scenario. These can easily be generalized when correlated uncertainties are present., we can define the GMM likelihood as
$L(Y|\vec{\theta})=\prod_{j=1}^{N_{\text{pt}}}\pi(y_{j},\Delta y_{j}|\vec{\theta})=\prod_{j=1}^{N_{\text{pt}}}\sum_{i=1}^{K}\omega_{i}\mathcal{N}(y_{j},\Delta y_{j}|\theta_{i}),$ (5)
@@
where the $i$’th base distribution $\mathcal{N}(y_{j},\Delta y_{j}|\theta_{i})$ is associated with a normalized weight $\omega_{i}$. Here $\theta_{i}$ corresponds to the fit parameters. For example, for the fits performed in the previous section $\theta_{i}=\left\{a_{4_{i}},a_{5_{i}}\right\}$ where $i=1,2,...K$. Therefore, performing a two parameter fit with K Gaussians, requires $2\times K$ parameters. To simplify the following expressions, we will abbreviate $\mathcal{N}$ and $\pi$ for the $i$’th base distribution and the $j$’th measurement as follows
@@
Notice that the number of original parameters $N_{\text{parm}}$ has increased in the GMM by a factor $K$. This means that we need to determine not only the central value of $K\times N_{\text{parm}}$ parameters but also the uncertainty on each of them. Later we will describe how this fit to $K\times N_{\text{parm}}$ parameters can be combined in a meaningful way to produced a fit and uncertainty on only $N_{\text{parm}}$ parameters.333 For example, there were two fit parameters ($a_{4}$ and $a_{5}$) in the LS fits in Sec. II, whereas in the GMM there would be $2\times K$ parameters. Ultimately, we want to be able determine the uncertainty on two parameters and not $2\times K$ parameters. In the GMM, the logarithm of the likelihood function is
@@
$-\ln L=-\sum_{j=1}^{N_{\text{pt}}}\ln\bigg{(}\mathcal{N}(y_{j},\Delta y_{j}|\theta)\bigg{)}=\frac{1}{2}\chi^{2}+\sum_{j=1}^{N_{\text{pt}}}\ln\Big{(}\sqrt{2\pi}\Delta y_{j}\Big{)}\ .$ (8)
@@ III.1 Performing parameter fits with the GMM
In order to determine the best fit parameters, we minimize the the log-likelihood defined in Eq. (7). We remark here that, conventinally the Expectation Maximization algorithm [31] is used to learn the best fit parameters of the mixture model. [^4] We remark here that it is possible to use the Expectation Maximization algorithm if we produce a large pseudo-data set and apply the Gaussian mixture model directly to the theory parameter space. The advantage of this algorithm is that it provides an iterative method that does not require information about the derivative of the likelihood. However, in our setup, a straightforward implementation of the Expectation Maximization algorithm is not possible since we are trying to fit a nonlinear function. For the purpose of demonstration, we use methods that rely on information about the derivative of the likelihood. Specifically, we use TensorFlow [32] to determine derivatives of the likelihood and minimize the log-likelihood. As described in the previous section, the number of parameters have increased and for $K$ Gaussians we have
$N_{\text{parm}}^{\text{GMM}}=K\times N_{\text{parm}}+K-1\ ,$ (9)
where, the last two terms, $K-1$, come from the constraint on the weights $\omega_{i}$ that are required to sum to unity. We carry out minimization of the likelihood by varying all $N_{\text{parm}}^{\text{GMM}}$ parameters. [^5] It is possible to determine the value of the weights $\omega_{i}$ in other more complicated ways since they indicate the degree of importance to each Gaussian. Here we take the approach that we learn the values of $\omega_{i}$ directly from the data and hence it is also varied during the minimization process. The gradient of the log-likelihood with respect to the fit parameters that belong to $i$-th Gaussian is
@@
$\frac{\partial}{\partial\theta_{i}}\ln L\bigg{|}_{\theta_{i}=\hat{\theta}_{i}}$ $=$ 00 0 (12)
@@
$\frac{\partial}{\partial w_{i}}\ln L\bigg{|}_{w_{i}=\hat{w}_{i}}$ $=$ $0$ (13)
In this work, we extremize the log-likelihood function by using the Adam optimizer with learning rate being $10^{-6}$ provided by the TensorFlow API [32]. Note that, in Eq. (10), the best-fit value of the prediction $y_{j}(\theta_{i})$ is weighted not only by the variance in measurements ($\Delta y_{i}$), but also by the weights $\omega_{i}$ and the factor $\mathcal{N}(y_{j}|\theta_{i})/\pi(y_{j},\Delta y_{j}|\vec{\theta})$ of the $i$-the base distribution, which becomes trivial for a single Gaussian. The Jacobian $\partial y_{j}(\theta_{i})/\partial\theta_{i}$, transforms information from the data space to the PDF space. The presence of this transformation makes it difficult to use the Expectation Maximum algorithm.
In this work we use the above equation as a constraint that determines the value of $\omega_{i}$. This is generally the approach taken for other unsupervised machine learning applications. In general, this strategy tends to assign higher weights to experiments with higher precision. If desired, it is possible to control the weights by increasing or decreasing the value of the uncertainties associated with a certain experimental data set. [^6] One method of trying to deal with inconsistent data sets is to artificially inflate uncertainties until we arrive at a consistent data set [16]. Although not described in this work, we remark that the GMM in conjunction with certain information criteria, provides a way to determine the extent to which it is necessary to inflate the uncertainty on data until the fits become consistent. Thus along with $\theta_{i}$, the weights $\omega_{i}$ are also varied in order to optimize the log-likelihood.
@@ III.2 Mean and Uncertainty in the GMM
It is straightforward to show that the expectation value of parameters is the weighted sum of the optimal parameters for individual base distributions,
@@
For the case of single Gaussian or when all the Gaussians have the same mean and covariance, the observed Fisher information matrix reduces to the usual Hessian matrix used to determine the covariance of the fit parameters in the LS method.777 For the GMM with multiple Gaussians, the likelihood function as defined in Eq. (7) includes all experimental points for each Gaussian. Thus, if the means of each Gaussian are identical, then this implies that the covariance, which is related to the second order derivative about the minimimum of the likelihood, must also be identical, c.f. Eq. (22). In general, the covariance matrices for Gaussians with identical means may be different in the GMM if the likelihoods were to be defined differently. Like the Hessian matrix, the information matrix can also be found with numerical methods. Here we use algorithmic differentiation [34] within TensorFlow to determine derivatives of the likelihood function. We note that the information matrix is only an estimate of the covariance of the GMM and it is possible to determine the covariance matrix as well, which will be described in detail elsewhere.
@@
×$\times\text{exp}\Big{[}-\frac{1}{2}\Big{(}(x-\theta_{i})^{T}\cdot\left(\text{cov}_{\text{GMM,i}}\right)^{-1}\cdot(x-\theta_{i})\Big{)}\Big{]}\ .$ (17)
@@
$-\frac{\partial^{2}}{\partial\theta\partial\theta}\ln L\bigg{|}_{\theta=\hat{\theta}}=\frac{1}{2}\frac{\partial^{2}}{\partial\theta\partial\theta}\chi^{2}\bigg{|}_{\theta=\hat{\theta}}=\sum_{j=1}^{N_{\text{pt}}}\frac{1}{\Delta y^{2}_{j}}\frac{\partial^{2}T(\theta)}{\partial\theta^{2}}\bigg{|}_{\theta=\hat{\theta}},$ (18)
@@
For the GMM, in order to provide a useful estimate of uncertainty, we need to reduce the $(KN_{\text{parm}})\times(KN_{\text{parm}})$ space to a $(N_{\text{parm}})\times(N_{\text{parm}})$ space in a meaningful way. To that end, let us consider the covariance of the GMM which is defined as
@@
Combining Eqs. (20) and (21), an estimate of the covariance of the GMM is given by
@@ IV Application of the GMM in a toy model of PDFs
In this section, we compare results obtained by the GMM method, defined in Eq. (7), to those obtained using the LS method. When estimating uncertainty using the LS method, we first diagonalize the Hessian and determine eigenvalue directions. We then determine values of the parameters that satisfy the $\Delta\chi^{2}=1$ condition and define the uncertainty on the fit parameters, which corresponds to the 68% confidence level interval. Here, we also implement the CT [1] and MSHT [2] tolerance criteria, which provide a more conservative estimate of uncertainty under the potential imperfections of fits. The comparison is presented in three cases of pseudo-data, including the two pseudo-data sets introduced in Sec. II.1.
@@ IV.1 Applying GMM to different scenarios
In this section we consider three cases. In each case we combine certain pseudo-data sets. We list and describe each of these cases below:
  • Case-1: Here we use pseudo-data sets #$\#~{}1$ and #$\#~{}2$ which were introduced in Sec. II.1. This case represents the scenario when there is strong tension between data sets. With this case, we will demonstrate how the GMM can provide an estimate of uncertainty in a statistically robust way.
  • Case-2: Here we combine pseudo-data sets #$\#~{}3$ and #$\#~{}4$. The central values used to generate these data sets are the same as those used in #$\#~{}1$ and #$\#~{}2$, however, their uncertainties are tripled by increasing the value of $\alpha$ in Eq. (2) by a factor of three, as shown in Table 3. The central values of data sets #$\#~{}1$ and #$\#~{}3$ as well as #$\#~{}2$ and #$\#~{}4$ do not overlap, because different random numbers ($r$), cf. Eq. (3), are used to generate them. With this combination we will be able to show how the GMM behaves when the errors of the data sets are large enough so that there is no tension. We can connect this to the alternative suggestions of dealing with uncertain data by artificially increasing the errors of the data sets [16].
  • Case-3: This case contains pseudo-data sets #$\#~{}5$ and #$\#~{}6$. The tension in this case is generated only via variation of the parameter $a_{2}$ The parameter $a_{2}$ controls the behaviour of the pseudo-PDF in the large-$x$ region, where the uncertainty is large. We fit this case with parameters $a_{2}$ and $a_{4}$. With this combination we demonstrate overfitting with the GMM method, and later in Sec. V.2 we will demonstrate how to deal with overfitting.
The true values of the parameters for cases 1 through 3 are presented in Table 3 and Fig. 3.
case pseudo-data $\alpha$ $N_{\text{pt}}$ $a_{0}$ $a_{1}$ $a_{2}$ $a_{3}$ $a_{4}$ $a_{5}$
Case-1 #1 0.05 50 30 0.5 2.4 4.3 2.4 -3.0
#2 0.05 50 30 0.5 2.4 4.3 2.6 -2.8
Case-2 #3 0.15 50 30 0.5 2.4 4.3 2.4 -3.0
#4 0.15 50 30 0.5 2.4 4.3 2.6 -2.8
Case-3 #5 0.05 50 30 0.5 2.4 4.3 2.4 -3.0
#6 0.05 50 30 0.5 1.4 4.3 2.4 -3.0
Table 3: Similar to Table 1, the summary of true values of the pseudo-PDF parameters and values of $\alpha$ in Eq. (2) for three cases. For comparison, we also include the true values of the pseudo-PDF parameters for Case-1 in this table.
@@
The best-fit obtained by the GMM method for the Case-1 is shown in Figs. 4 and 4 and Table 4. For the GMM likelihood function, we choose to fit with a mixture of two Gaussian distributions. As a result, the mean values of each Gaussian, can describe the two distinct pseudo-data sets well and in fact are very similar to what we obtain for SG-A and SG-B. We find the weights $\omega_{i}$ to be roughly equal, which is a result of the errors in each of the data sets #$\#1$ and #$\#2$ being almost equal and also having the same number of data points. In Figs. 4and 4, the pseudo-PDF uncertainty estimated by the GMM is compared to uncertainties predicted using the LS method. For $0.110^{-3}<x<$ 0.1, the (small) error band of SG-C pseudo-PDF exclude those of SG-A and SG-B fits, the error band estimated by the GMM is consistent and includes areas covered by both SG-A and SG-B pseudo-PDFs. For $x>0.1$, where the error in the data sets are large, uncertainties from SG-C and the GMM overlap. Fig. 4 shows the results of the different fits in the fit parameter space. The 1-$\sigma$ correlation ellipse estimated by SG-C is small, as has been mentioned before, and does not overlap with SG-A and SG-B replicas. On the other hand, the GMM correlation ellipse is large enough and, importantly, also possesses a different axes orientations, that enables the GMM covariance to overlap with some of the replicas of both SG-A and SG-B. Note, that we have represented the GMM by a single covariance matrix but this should not be confused with the covariance arising from a single Gaussian, since the shape of the GMM likelihood is multi-modal and not uni-modal.
For Case-2, shown in Fig. 5, the tension in the data set has been reduced by increasing the error on the data, so that SG-C describes the combined pseudo-data set with $\chi^{2}_{\#3+\#4}/N_{\text{pt}}=1.04$. Even for such a mild tension, the fits SG-A and SG-B do not overlap with SG-C. On the other hand, the GMM fit overlaps with both SG-A and SG-B. The difference between the axes orientations of the correlation ellipses between SG-C and the GMM is small. Although, the GMM method produces a larger uncertainty which is more consistent with SG-A and SG-B, the differences with SG-C are small and we need a way to determine whether we should stick with SG-C or with the GMM method. We will return to a discussion of this in Sec. V.2.

### v2

@@ II.1 Least Squares Fits and Tension
Figure 2: Panel (a): the pseudo-PDFs determined from each of the fits are shown. The dot-dashed lines refer to the central PDFs for SG-A (green), SG-B (yellow) and the combination SG-C (blue). Uncertainties at 1 standard deviation in the replica sets are shown by the dotted lines for each fit. For the LS fit C, the blue shaded region represents the Hessian uncertainty with $\Delta\chi^{2}=1$, which deviates from the Monte Carlo uncertainty by a negligible amount. The sizes of uncertainties are small with respect to the magnitude of the pseudo-PDFs. Panel (b): similar to panel (a), but for ratios to the central value of the Hessian set of the SG-C fit. Panel (c): the values of parameters $a_{4}$ and $a_{5}$ in replica sets for SG-A, SG-B and SG-C are shown along with the 1 and 3 standard deviation uncertainties for SG-C given by the solid curves.
Before we conclude this section, we comment on alternate methods to deal with such inconsistencies in fits. For example, as described in Appendix C, the Particle Data Group (PDG) [38], recommends increasing the uncertainty on data sets by a scale factor $S=\sqrt{\chi^{2}/(N_{\text{pt}}-1)}$. For SG-C, this implies scaling the uncertainty on all the data by $S\simeq 1.56$, which, for our simple example, translates to setting a tolerance $T=S\simeq 1.56$. Increasing the error with the scale factor, or equivalently the tolerance, does not shift the central value of the fit and only increases the uncertainty marginally. In Sec. IV.2, we explore even larger values of the tolerance and we will see that such a small value of $T\simeq 1.56$ is not sufficient to reproduce uncertainty bands that represent the tension between the data sets.
It is crucial to understand why the uncertainty band of SG-C, even with the increased tolerance ($T=S\simeq 1.56$), is unsatisfactory. It is easy to imagine, that the true theory value could easily be close to the central values of SG-A or SG-B, i.e. there is a large range of possible values for the true theory parameters, given the data. On the other hand, the central values of SG-A and SG-B lie well outside the uncertainty band of the combined LS fit SG-C, even with $T\simeq 1.56$. Hence, the uncertainty bands of SG-C, do not provide an accurate representation of the situation. In other words, precision is lost and the uncertainties must be increased in the region of tension in order to avoid biasing the fits and to accommodate our lack of information about the source of tension. We describe a simpler example using $W$-boson mass in Appendix C, where we show how a naive combination of data with tension can bias the fits even after increasing the scale factor $S$.
The method using the GMM, to be presented in the next section, provides an alternate way to combine inconsistent data sets with large tension and provides a more accurate determination of the uncertainties. Furthermore, if desired, it can also provide an alternate estimate for the PDG scale factor $S$ required to make data sets consistent. We describe some of these ideas briefly in Sec. V.2 and Appendix C.
@@ III Gaussian mixture model in PDF fit
As mentioned in Sec. II, the likelihood from each of the data sets (pseudo-data #1 and pseudo-data #2) are significantly different and need to be combined in a meaningful way. We propose the use of the Gaussian Mixture Model to construct the combined likelihood for data with tension.
The GMM was first used by Karl Pearson in 1894 [39]. It is now used extensively as an unsupervised machine learning technique to classify data [25]. Here, we use it to determine the uncertainty estimate on theoretical fit parameters given a set of discrepant data. In Appendix B, we present the more intuitive Bayesian description of the GMM. Here, we present the frequentist approach by starting with an ansatz for the likelihood function. In general, the GMM is a mixture of $K$ base Gaussian distributions. Given the data set $Y=\{(y_{1},\Delta y_{1}),(y_{2},\Delta y_{2}),\cdots,(y_{N_{\text{pt}}},\Delta y_{N_{\text{pt}}})\}$, where $\Delta y_{j}$ is the uncertainty associated with the measured value $y_{j}$,444 For clarity of presentation, we use an uncorrelated data set and present formulas for this scenario. These can easily be generalized when correlated uncertainties are present. we can define the GMM likelihood as
$L(Y|\vec{\theta})$ $=$ $\prod_{j=1}^{N_{\text{pt}}}\pi(y_{j},\Delta y_{j}|\vec{\theta})=\prod_{j=1}^{N_{\text{pt}}}\left(\sum_{i=1}^{K}\omega_{i}\mathcal{N}(y_{j},\Delta y_{j}|\theta_{i})\right)\ ,$
@@
where the $i$’th base distribution $\mathcal{N}(y_{j},\Delta y_{j}|\theta_{i})$ is associated with a normalized weight $\omega_{i}$.
Here, the index $j$ labels each experimental data point, index $i$ labels each Gaussian with theory parameters $\theta_{i}$ that are estimated by maximizing the likelihood above. In the example fit that we use $\theta_{i}=\left\{a_{4_{i}},a_{5_{i}}\right\}$ where i=1,2,Ki=1,2,...K. For example, if $K=2$, then $\theta_{1}=\left\{a_{4_{1}},a_{5_{1}}\right\}$ and $\theta_{2}=\left\{a_{4_{2}},a_{5_{2}}\right\}$ and we have effectively introduced four parameters to fit. If $K=1$, then $\theta_{1}=\left\{a_{4_{1}},a_{5_{1}}\right\}=\left\{a_{4},a_{5}\right\}$ and the log likelihood reduces to the usual $\chi^{2}$ likelihood used in Sec. II.1. For $K>1$, in addition to the (K1)(K-1) independent weight parameters $w_{i}$, we have $N_{\text{parm}}\times K$ parameters that are varied to maximize the likelihood, where $N_{\text{parm}}$ is the number of parameters in the fitting function $g(x)$ that are being estimated. Since the number of original parameters $N_{\text{parm}}$ has increased in the GMM by a factor $K$, we need to determine not only the central value of $K\times N_{\text{parm}}$ parameters but also the uncertainty on each of them. Later, we will describe how this fit to $K\times N_{\text{parm}}$ parameters can be combined in a meaningful way to produce a fit and uncertainty on only $N_{\text{parm}}$ parameters.555 For example, there were two fit parameters ($a_{4}$ and $a_{5}$) in the LS fits in Sec. II, whereas in the GMM there would be $2\times K$ parameters. Ultimately, we want to be able determine the uncertainty on two parameters and not $2\times K$ parameters.
To simplify the following expressions, we will abbreviate $\mathcal{N}$ and $\pi$ for the $i$’th base distribution and the $j$’th measurement as follows
@@
In the GMM, the logarithm of the likelihood function is
@@
$\ln L=-\sum_{j=1}^{N_{\text{pt}}}\ln\bigg{(}\mathcal{N}(y_{j},\Delta y_{j}|\theta)\bigg{)}=\frac{1}{2}\chi^{2}+\sum_{j=1}^{N_{\text{pt}}}\ln\Big{(}\sqrt{2\pi}\Delta y_{j}\Big{)}\ .$ (6)
@@
See Appendix B.2, for a more intuitive explanation of this idea using a Bayesian formalism.
@@ III.1 Performing parameter fits with the GMM
In order to determine the best fit parameters, we minimize the the log-likelihood defined in Eq. (5). We remark here that, conventionally the Expectation Maximization algorithm [25] is used to learn the best fit parameters of the mixture model. [^6] We remark here that it is possible to use the Expectation Maximization algorithm if we produce a large pseudo-data set and apply the Gaussian mixture model directly to the theory parameter space. The advantage of this algorithm is that it provides an iterative method that does not require information about the derivative of the likelihood. For the purpose of demonstration, we use methods that rely on information about the derivative of the likelihood. Specifically, we use TensorFlow [40] to determine derivatives of the likelihood and minimize the log-likelihood. As described in the previous section, the number of parameters have increased and for $K$ Gaussians we have
$N_{\text{parm}}^{\text{GMM}}=K\times N_{\text{parm}}+(K-1)\ ,$ (7)
where, the last term, K1K-1, come from the constraint on the weights $\omega_{i}$ that are required to sum to unity. We carry out minimization of the likelihood by varying all $N_{\text{parm}}^{\text{GMM}}$ parameters. [^7] It is possible to determine the value of the weights $\omega_{i}$ in other more complicated ways since they indicate the degree of importance to each Gaussian. Here we take the approach that we learn the values of $\omega_{i}$ directly from the data and hence it is also varied during the minimization process. The gradient of the log-likelihood with respect to the fit parameters that belong to $i$-th Gaussian is
@@
$\frac{\partial}{\partial\theta_{i}}\ln L\bigg{|}_{\theta_{i}=\hat{\theta}_{i}}$ $=$ $0$ (10)
@@
$\frac{\partial}{\partial w_{i}}\ln L\bigg{|}_{w_{i}=\hat{w}_{i}}$ $=$ $0.$ (11)
In this work, we extremize the log-likelihood function by using the Adam optimizer with learning rate being $10^{-6}$ provided by the TensorFlow API [40]. Note that, in Eq. (8), the best-fit value of the prediction $y_{j}(\theta_{i})$ is weighted not only by the variance in measurements ($\Delta y_{i}$), but also by the weights $\omega_{i}$ and the factor $\mathcal{N}(y_{j}|\theta_{i})/\pi(y_{j},\Delta y_{j}|\vec{\theta})$ of the $i$-the base distribution, which becomes trivial for a single Gaussian. The Jacobian $\partial y_{j}(\theta_{i})/\partial\theta_{i}$, transforms information from the data space to the PDF space. The presence of this transformation makes it difficult to use the Expectation Maximum algorithm directly.
In this work we use the above equation as a constraint that determines the value of $\omega_{i}$. This is generally the approach taken for other unsupervised machine learning applications. In general, this strategy tends to assign higher weights to experiments with higher precision. If desired, it is possible to control the weights by increasing or decreasing the value of the uncertainties associated with a certain experimental data set. [^8] One method of method of managing inconsistent data sets is to artificially inflate uncertainties until we arrive at a consistent data set [16]. In Appendix C, we describe how the GMM, in conjunction with certain information criteria, provides a way to determine the extent to which it is necessary to inflate the uncertainty on data until the fits become consistent. Thus along with $\theta_{i}$, the weights $\omega_{i}$ are also varied in order to optimize the log-likelihood.
@@ III.2 Mean and Uncertainty in the GMM
It is straightforward to show that the expectation value of parameters is the weighted sum of the optimal parameters for individual base distributions [42],
@@
Here $\hat{\theta}_{i}$ are the estimated values of the $N_{\text{parm}}$ parameters of the fitted function $g(x)$ and there are $K$ such $\hat{\theta}_{i}$.
@@
When using a single Gaussian or when all the Gaussians have the same mean and covariance, the observed Fisher information matrix reduces to the usual Hessian matrix used to determine the covariance of the fit parameters in the LS method.999 For the GMM with multiple Gaussians, the likelihood function as defined in Eq. (5) includes all experimental points for each Gaussian. Thus, if the means of each Gaussian are identical, then this implies that the covariance, which is related to the second order derivative about the minimum of the likelihood, must also be identical, c.f. Eq. (20). In general, the covariance matrices for Gaussians with identical means may be different in the GMM if the likelihoods were to be defined differently. Like the Hessian matrix, the information matrix can also be found with numerical methods. Here we use algorithmic differentiation [43] within TensorFlow to determine derivatives of the likelihood function. We note that the information matrix is only an estimate of the covariance of the GMM and it is possible to determine the covariance matrix as well, which will be described in detail elsewhere.
@@
$\times\text{exp}\Big{[}-\frac{1}{2}\Big{(}(x-\theta_{i})^{T}\cdot\left(\text{cov}_{\text{GMM,i}}\right)^{-1}\cdot(x-\theta_{i})\Big{)}\Big{]}\ .$ (15)
@@
$\frac{\partial^{2}}{\partial\theta\partial\theta}\ln L\bigg{|}_{\theta=\hat{\theta}}=\frac{1}{2}\frac{\partial^{2}}{\partial\theta\partial\theta}\chi^{2}\bigg{|}_{\theta=\hat{\theta}}=\sum_{j=1}^{N_{\text{pt}}}\frac{1}{\Delta y^{2}_{j}}\frac{\partial^{2}T(\theta)}{\partial\theta^{2}}\bigg{|}_{\theta=\hat{\theta}}$, (16)
@@
For the GMM, in order to provide a useful estimate of uncertainty, we need to reduce the $(KN_{\text{parm}})\times(KN_{\text{parm}})$ space to a $(N_{\text{parm}})\times(N_{\text{parm}})$ space in a meaningful way. To that end, let us consider the covariance of the GMM which is defined as [42]
@@
Combining Eqs. (18) and (19), an estimate of the covariance of the GMM is given by [42]
@@ IV Application of the GMM in a toy model of PDFs
In this section, we compare results obtained by the GMM method, using the likelihood defined in Eq. (5), to those obtained using the LS method. When estimating uncertainty using the LS method, we first diagonalize the Hessian and determine eigenvalue directions. We then determine values of the parameters that satisfy the $\Delta\chi^{2}=1$ condition and define the uncertainty on the fit parameters, which corresponds to the 68% confidence level interval. Here, we also implement the CTEQ-TEA [1] and MSHT [2] tolerance criteria, which provide a more conservative estimate of uncertainty under the potential imperfections of fits.
Below we present the results of three different combinations of pseudo-data sets, including also the two pseudo-data sets that were discussed in Sec. II.
@@ IV.1 Applying GMM to different scenarios
In this section we consider three cases of combining different pairs of pseudo-data sets. Details of how each of these psuedo-data sets are generated can be found in Appendix A.
Here, we list and describe each of these cases below:
  • Case-1: Here we use pseudo-data sets #1\#~1 and #2\#~2 which were introduced in Sec. II. The pseudo-data sets are plotted in Fig. 1.This case represents the scenario when there is strong tension between data sets. With this case, we will demonstrate how the GMM can provide a more accurate estimate of uncertainty.
  • Case-2:
    Here we combine pseudo-data sets #3\#~3 and #4\#~4, which are shown in Fig. 3(a) and Fig. 3(b). These two data sets are similar to data sets #1\#~1 and #2\#~2, except that their uncertainties are tripled and thus have less tension.
    With this combination we will be able to show how the GMM behaves when the errors of the data sets are large enough so that there is less tension. We can connect this to the alternative suggestions of dealing with uncertain data by artificially increasing the errors of the data sets [16].
  • Case-3:
    This case contains pseudo-data sets #5\#~5 and #6\#~6, which are shown in Fig. 3(c) and Fig. 3(d). The tension between these two pseudo-data sets resides in the large-$x$ region
    where the uncertainty is large. We fit this case with parameters $a_{2}$ and $a_{4}$. With this combination we demonstrate overfitting with the GMM method, and later in Sec. V.2 we demonstrate how to address overfitting.
@@
The best-fit obtained by the GMM method for Case-1 is shown in Figs. 4 and 4 and in Table 2. We choose to fit the GMM likelihood function with a mixture of two Gaussian distributions. As a result, the mean values of each Gaussian, can describe the two distinct pseudo-data sets well and are in agreement with the fits SG-A and SG-B. We find the weights $\omega_{i}$ to be roughly equal, which is a result of the errors in each of the data sets #1\#1 and #2\#2 being almost equal and also having the same number of data points. In Figs. 4and 4, the pseudo-PDF uncertainty estimated by the GMM is compared to uncertainties predicted using the LS method. For $10^{-3}<x<0.1$, the (small) error band of SG-C pseudo-PDF exclude those of SG-A and SG-B fits, the error band estimated by the GMM is consistent and includes areas covered by both SG-A and SG-B pseudo-PDFs. For $x>0.1$, where the error in the data sets are large, uncertainties from SG-C and the GMM overlap. Fig. 4 shows the results of the different fits in the fit parameter space. The 1-$\sigma$ correlation ellipse estimated by SG-C is small, as has been mentioned before, and does not overlap with SG-A and SG-B replicas. On the other hand, the GMM correlation ellipse is large enough and, importantly, also possesses a different axes orientations, that enables the GMM covariance to overlap with some of the replicas of both SG-A and SG-B. The essential feature we see here is that the tension in the pseudo-data sets lies in the region $10^{-2}<x<10^{-1}$ (see for example Fig. 4). The GMM expands errors in this region, whereas the $\chi^{2}$ method with tolerance displays a smaller uncertainty band in this region and expands the error in the region $x>10^{-1}$. The GMM therefore provides a better estimate of the probability distribution of the theory parameters given the discrepancies in the data. Note that we have represented the GMM by a single covariance matrix but this should not be confused with the covariance arising from a single Gaussian, since the shape of the GMM likelihood is multi-modal and not uni-modal.
For Case-2, shown in Fig. 5, the tension in the data set has been reduced by increasing the error on the data, so that SG-C describes the combined pseudo-data set with $\chi^{2}_{\#3+\#4}/N_{\text{pt}}=1.04$. Even for such a mild tension, the fits SG-A and SG-B do not overlap with SG-C. On the other hand, the GMM fit overlaps with both SG-A and SG-B. The difference between the axes orientations of the correlation ellipses between SG-C and the GMM is small. Although, the GMM method produces a larger uncertainty which is more consistent with SG-A and SG-B, the differences with SG-C are small and we need a way to determine whether we should use SG-C or the GMM method. We will return to a discussion of this in Sec. V.2.